### v1

# Jiahao Huo1,3, Yibo Yan1,2, Boren Hu1,2, Yutao Yue1,2, Xuming Hu1,2

Abstract
Projecting visual features into word embedding space has become a significant fusion strategy adopted by Multimodal Large Language Models (MLLMs). However, its internal mechanisms have yet to be explored. Inspired by multilingual research, we identify domain-specific neurons in multimodal large language models. Specifically, we investigate the distribution of domain-specific neurons and the mechanism of how MLLMs process features from diverse domains. Furthermore, we propose a three-stage framework for language model modules in MLLMs when handling projected image features, and verify this hypothesis using logit lens. Extensive experiments indicate that while current MLLMs exhibit Visual Question Answering (VQA) capability, they may not fully utilize domain-specific information. Manipulating domain-specific neurons properly will result in a 10% change of accuracy at most, shedding light on the development of cross-domain, all-encompassing MLLMs in the future. Our code will be released upon paper notification.
[Uncaptioned image]MMNeuron: Discovering Neuron-Level Domain-Specific Interpretation in Multimodal Large Language Model
Jiahao Huo1,3, Yibo Yan1,2, Boren Hu1,2, Yutao Yue1,2, Xuming Hu1,2thanks:  Corresponding Author 1 The Hong Kong University of Science and Technology (Guangzhou) 2 The Hong Kong University of Science and Technology, 3 Tongji University {jiahaohuotj, yanyibo70, huboren99}@gmail.com, {yutaoyue, xuminghu97}@hkust-gz.edu.cn
1 Introduction
Neuron Analysis, which interprets activation of neurons as the recall of learned knowledge in deep neural networks, has been widely adopted by researchers to understand the inner workings of models Sajjad et al. (2022); Fan et al. (2024). Prior studies have confirmed that certain neurons within deep neural networks play important roles in learning particular concepts Oikarinen and Weng (2022); Bai et al. (2024); Xiao et al. (2024), preserving factual knowledge Chen et al. (2024); Dai et al. (2022); Niu et al. (2024) as well as solving specific tasks Stanczak et al. (2022). Beyond enhancing model interpretability, current practical applications of Neuron Analysis include model distillation Dalvi et al. (2020), knowledge editing Chavhan et al. (2024); Pan et al. (2023), and controllable generation Bau et al. (2019); Kojima et al. (2024). Central to such endeavors is the identification of neurons responsible for target scenarios.
Figure 1: Neuron analysis in previous language-specific setting of large language model (a) and our domain-specific setting of multimodal large language model (b).
As illustrated in Figure 1 (a), recent studies have focused on interpreting the multilingual capabilities of pre-trained large language models (LLMs) under the view of language-specific neurons, which are neurons uniquely responsible for particular languages. For instance, Kojima et al. (2024) identified such neurons in pre-trained decoder-based language models, demonstrating that tampering with a few language-specific neurons significantly alters the occurrence probability of target language in text generation. Similarly, Zhao et al. (2024c) detected language-specific neurons by measuring the significance of neurons when processing multilingual inputs and proposed a workflow of LLMs handling multilingual tasks. Moreover, Tang et al. (2024) used language activation probability entropy (LAPE) to identify language-specific neurons, demonstrating that activating or deactivating certain neurons can change the language of the model’s output. On the other hand, it has also been confirmed that neurons in text-only transformers can understand visual features extracted by a vision encoder Schwettmann et al. (2023).
These findings have prompted an interesting question: Do similar mechanisms exist in multimodal large language models (MLLMs) during the processing of features from different visual domains? As shown in Figure 1(b), we aim to apply the mechanism similar to multilingual neuron analysis  Tang et al. (2024) to current representative open-source MLLMs, including LLaVA-NeXT Liu et al. (2024a) and InstructBLIP Dai et al. (2024). The aforementioned models extract image features via a pre-trained vision encoder and project these features into the word embedding space. These post-projection visual features are concatenated with language features and fed into the model’s LLM module to generate text outputs.
Figure 2: PCA visualization of image embeddings extracted through CLIP’s image encoder.
Specifically, we investigate the activation patterns of neurons in MLLMs’ feed-forward network (FFN) layers across corpora from five distinct domains, identifying less than 1% as domain-specific neurons. The datasets we utilized include LingoQA Marcu et al. (2023), RS-VQA (HR) Lobry et al. (2020), PMC-VQA Zhang et al. (2023b), DocVQA Mathew et al. (2021) and VQAv2 Goyal et al. (2017), covering domains such as Auto Driving, Remote Sensing, Medicine, Document, and Common Scenes. Figure 2 highlights the clustering and separation of image features across the domains. Image examples of these domains can also be found in Appendix A. Based on our experiment results, we argue that differences exist among these visual domains and that the vision encoder and LLM modules in MLLMs exhibit distinct patterns for these domains. Furthermore, we propose a three-stage framework based on the distribution of domain-specific neurons among MLLM’s LLM layers, where post-projection visual features are processed by LLM. To validate our hypothesis, we employ logit lens nostalgebraist (2020) to decode the hidden states of LLM’s intermediate layers to visualize the feature transformation within transformer models Vaswani et al. (2017).
Our main contributions are as follows:
  • We identify the presence of domain-specific neurons in representative MLLMs, which is vital for interpreting domain-specific features.
  • We analyze the impact of domain-specific neurons, indicating that both LLaVA-NeXT and InstructBLIP do not fully utilize domain-specific information in particular domains.
  • We compare features from various domains through the lens of domain-specific neurons, revealing that images from different domains vary in conceptual depth.
  • We propose a three-stage framework of language models in MLLMs when processing projected image features, shedding light on the internal mechanisms by which image features align with word embeddings.
To the best of our knowledge, we are the first to investigate domain-specific neurons in the multimodal field, although there are already insightful discussions on visual representations in MLLMs Schwettmann et al. (2023); Zhao et al. (2024a). Our findings can reveal the neuron-level similarity and distinction among these domains, offering insights to understand and enhance the cross-domain potential of current MLLMs.
2 Related Work
2.1 Neuron Analysis
Neuron analysis has been recently widely explored in computer vision and natural language processing, which views neuron activation as the recall of learned knowledge Mu and Andreas (2020); Sajjad et al. (2022)Bau et al. (2017) propose to automatically inspect the functionality of each visual neuron in CNNs by evaluating the alignment between individual hidden units. Hernandez et al. (2021); Oikarinen and Weng (2022); Bai et al. (2024) further extend this method to open-ended by labeling hidden neurons in visual models with natural language descriptions. Neuron analysis has also been adopted to analyze language models, including the ability of sentiment analysis Radford et al. (2017), machine translation Mu et al. (2024), knowledge storing Dai et al. (2022); Zhao et al. (2024b); Chen et al. (2024) and task solving Wang et al. (2022). Recent research has associated specific neurons in LLMs with their multilingual ability, describing these neurons as language-specific neurons Tang et al. (2024); Zhao et al. (2024c). Inspired by their work, we further expand this idea to the multimodal domain, being the first to analyze the domain-specific neurons in MLLMs.
2.2 Visual Representation in Word Embedding Space
Aligning image features within the word embedding space of LLMs has been one of the dominant frameworks adopted by current open-source MLLMs. Large Language and Vision Assistant (LLaVA) and its variants Liu et al. (2024b, 2023a, a) use a simple linear layer to connect image features extracted by the vision encoder of CLIP Radford et al. (2021) into the word embedding space of LLMs Touvron et al. (2023); Chiang et al. (2023); Jiang et al. (2023). Instead of concatenating post-projected embeddings directly with language instructions, InstructBLIP Dai et al. (2024) employs a Q-Former to extract image features based on the instruction, which was more efficient. Similarly, MiniGPT-4 Zhu et al. (2023) gained image features through pre-trained ViT Dosovitskiy et al. (2020) or Q-Former Li et al. (2022), which are then projected into the word space by a linear layer. Although such a framework has gained remarkable performance in various multimodal tasks Antol et al. (2015); Chen et al. (2015); Liu et al. (2023b), the mechanism through which image tokens are processed by the LLM module still needed to be clarified. Our research has shed light on the interpretation of how MLLM understands the image tokens.
2.3 Cross-domain MLLM
Researchers have managed to fine-tune current general-domain MLLMs on specific domain corpus. For example, Kuckreja et al. (2024) train MLLM on the Remote Sensing multimodal dataset using LLaVA-1.5 architecture. LLaVA-Med Li et al. (2023) was initialized with the general-domain LLaVA and then continuously trained in a curriculum learning fashion, while VLAAD Park et al. (2024) opts for Video-LLaMA Zhang et al. (2023a) as the foundational model to assist LLM in comprehending video data from auto driving scenarios. There are also researches trying to enhance MLLM’s performance in specific domains  Bazi et al. (2024); Seyfioglu et al. (2023); Shao et al. (2023); Tian et al. (2024). Despite these efforts, it has also been proved that general-domain MLLMs without further domain-specific fine-tuning have demonstrated some cross-domain capability on some less common domains Verma et al. (2024); Lu et al. (2023). In our research, we select virgin (i.e., without further fine-tuning) LLaVA-NeXT and InstructBLIP as our baseline, hoping to bring insights into the interpretation of general-domain MLLM’s cross-domain potential and the development of all-around MLLMs qualified for different domains.
3 Method
3.1 Neuron Activation Detection
Figure 3: The overall framework of our proposed MMNeuron method (taking LLaVA architecture as an example), which can be applied to any MLP layers with an activation layer in multimodal large language models.
A prevalent framework for vision-language models involves utilizing a pre-trained vision encoder to extract image features $Z_{v}$ from image $X_{v}$. These features are then aligned with the word embedding space via a projection module, yielding post-projection features denoted as $H_{v}$. This process can be formalized as follows:
$H_{v}=f_{\Pi}(Z_{v}),\quad\text{with}\quad Z_{v}=f_{\Theta}(X_{v}).$ (1)
Here, $f_{\Pi}(\cdot)$ and $f_{\Theta}(\cdot)$ represent the projection module parameterized by $\Pi$ and the vision encoder parameterized by $\Theta$. In LLaVA, the projection module is a simple linear layer, whereas in InstructBLIP, it is implemented via a Q-Former Li et al. (2022). The post-projection features are then concatenated with language instruction embeddings $H_{q}$ and fed into an LLM to generate text answer $X_{a}$:
$X_{a}=f_{\Phi}([H_{v},H_{q}]),$ (2)
where $f_{\Phi}(\cdot)$ refers to the language model parameterized by $\Phi$.
For each Feed-Forward Network (FFN) layer, we consider every activation function as a neuron, as depicted in Figure 3. Given the hidden state $h^{i}\in\mathbb{R}^{d}$ of the input of the $i$-th FFN layer, the output of the FFN layer can be expressed as:
$h^{i+1}={\rm act\_fn}(h^{i}W_{1}^{i})W_{2}^{i},$ (3)
where ${\rm act\_fn}(\cdot)$ denotes the activation function (e.g., GELU in Figure 3), and $W_{1}^{i}\in\mathbb{R}^{d\times s}$ and $W_{2}^{i}\in\mathbb{R}^{s\times d}$ represent the parameters of first Linear Layer and second Linear Layer. Here, $s$ is the intermediate size of FFN layer. Therefore, there are $s$ neurons in this FFN layer. Conventionally, the $j$-th neuron inside the $i$-th FFN layer is activated only if its respective activation value $act\_fn(h^{i}W_{1}^{i})_{j}$ exceeds zero Nair and Hinton (2010).
3.2 Domain-Specific Neuron Selection
Our selection method is based on Tang et al. (2024). For each domain $D_{i},i=1,2,...,k$, we feed its image-text corpus into MLLM, and record the activated frequency of each neuron $u$ as well as the total token nums $N_{u,i}$ [^1] Note that neurons in the vision encoder and language model may receive different numbers of tokens, since projected image features are concatenated with language embeddings before being fed into language model.. The activation probability of a neuron $u$ in domain $D_{i}$ is denoted as:
$p_{u,i}=\frac{M_{u,i}}{N_{u,i}}.$ (4)
We denote the probability distribution of neuron $u$ across all domains as $P_{u}$:
$P_{u}=(p_{u,1},p_{u,2},...,p_{u,k}).$ (5)
The distribution can be normalized to a valid probability distribution through L1 normalization:
$\begin{split}P_{u}^{\prime}=(p_{u,1}^{\prime},p_{u,2}^{\prime},...,p_{u,k}^{\prime}),\\ \text{where}\quad P_{u,i}^{\prime}=\frac{P_{u,i}}{\sum_{j=1}^{k}P_{u,j}}.\end{split}$ (6)
Such a valid probability distribution allows us to calculate its corresponding entropy, termed domain activation probability entropy (DAPE):
$DAPE_{u}=-\sum_{j=1}^{k}p_{u,j}\log{p_{u,j}}.$ (7)
Intuitively, a lower entropy indicates a tendency for activation in response to one or two domains, with reduced activation probabilities for others. Thus, neurons with low DAPE are designated as domain-specific neurons, following Tang et al. (2024). In our work, we select those neurons with the bottom 1% DAPE scores as domain-specific neurons.
Upon identifying domain-specific neurons, we further analyze their specificity across five domains. A domain-specific neuron $u$ is considered specific to domain $D_{j}$ if its activation probability $p_{u,j}$ exceeds a predefined threshold.
3.3 Latent Embeddings Interpretation
Figure 4: General Framework of logit len analysis, where it takes the hidden state at an intermediate layer (e.g., $h1$ above), and convert the hidden state into logits with the unembedding layer. Note that Emb, Pos Emb, Res, and Unemb stand for Embedding, Position Embedding, Residual Layer, and Unembedding, respectively.
Consider a transformer model, where its $l$-th layer updates the representation as follows:
$h_{l+1}=h_{l}+F_{l}(h_{l}).$ (8)
Here, $F_{i}$ is the residual output of layer $i$. By applying Equation 8 recursively, the final output logits of model can be written as a function of an arbitrary hidden state $h_{i}$ at the $i$-th layer:
$logit(h_{l})=LayerNorm(h_{l}+\sum_{i=l}^{L}F_{i}(h_{i}))W_{U},$ (9)
where the term $\sum_{i=l}^{L}F_{i}(h_{i})$ represents the residual updates in the subsequent layers, and $W_{U}$ denotes the so-called unembedding matrix. The logit lens approach involves setting the residuals to zero Belrose et al. (2023):
$LogitLens(h_{l})=LayerNorm(h_{l})W_{U}.$ (10)
As shown in Figure 4, the logit lens decodes the hidden states of the transformer’s intermediate layers into the distribution over the vocabulary, which can be used to interpret the model’s latent embeddings nostalgebraist (2020). Ideally, the decoded distribution converges monotonically toward the next token predicted by the model.
We apply this trick to decode the hidden states of the language model, which allows us to understand the transformation of post-projection features within the language model module of the MLLM.
4 Experiment
In this section, we present empirical evaluation to elucidate the impact of domain-specific neurons, showing the potential mechanism of how MLLMs interpret image and language instructions.
4.1 Experimental Setup
4.1.1 Models
We study two public models: LLaVA-NeXT Liu et al. (2024a) and InstructBLIP Dai et al. (2024). The former utilizes a simple MLP layer to project image features extracted by CLIP’s vision encoder into the word embedding space. The latter, however, employs the Q-Former Li et al. (2022) to refine the image features extracted by ViT Dosovitskiy et al. (2020). Specifically, we select llava-v1.6-vicuna-7b-hf [^2] https://huggingface.co/llava-hf/llava-v1.6-vicuna-7b-hf and Instructblip-vicuna-7b [^3] https://huggingface.co/Salesforce/instructblip-vicuna-7b, both of which use Vicuna-7b Chiang et al. (2023) as the language model base. The number of neurons in llava-v1.6-vicuna-7b-hf and Instructblip-vicuna-7b are 454.7K and 665.6K, respectively.
4.1.2 Dataset and Metrics
We select five datasets representing five different domains, namely, VQAv2 Goyal et al. (2017) for common scenes, PMC-VQA Zhang et al. (2023b) for Medical domain, DocVQA Mathew et al. (2021) for Document domain, LingoQA Marcu et al. (2023) for Auto Driving domain and RS-VQA Lobry et al. (2020) for Remote Sensing domain. For LingoQA, visual instruction for each question includes multiple images, as shown in Figure 13(a). We prepare image-question pairs of nearly the same token numbers for each domain during identifying, around 20 million tokens in LLaVA-NeXT. During evaluation, the scale of the validation set is aligned with LingoQA to make a fair comparison. For DocVQA, we report Average Normalized Levenshtein Similarity (ANLS) score Biten et al. (2019) followed by the official benchmark. For LingoQA, we use the score of Lingo-Judge Marcu et al. (2023) with the official implementation. For all other datasets, we report the top-1 accuracy (%) as the metric.
4.1.3 Implementation Details
We adhere to the default prompt templates from the official repository or the original paper during evaluation, with an additional role description for the auto-driving scenes. For more details, please refer to Appendix B. We perform the forward pass without padding or truncation during the identification process. When evaluating models across different datasets, we employ beam search with max_length of 512 and num_beams of 5 to generate answers. The temperature and length_penalty arguments are set as 0.9 and -1, respectively.
4.2 Results & Discussion
4.2.1 Distribution of Domain-specific Neurons
(a) Distribution of domain-specific neurons in InstructBLIP.
(b) Distribution of domain-specific neurons in LLaVA-NeXT. $\star$: The MLP projector of LLaVA-NeXT consists of only one single layer.
Figure 5: Layer-wise Distribution of domain-specific neurons in different modules.
Baseline Module VQAv2 PMC-VQA LingoQA DocVQA RS-VQA
LLaVA-NeXT Vision Encoder 65 233 168 409 465
MLP Projector 8 13 13 11 20
LLM 683 915 1536 423 2120
InstructBLIP Vision Encoder 94 488 279 916 891
Q-Former 39 206 334 175 72
LLM 410 774 1567 556 1419
Table 1: The number of neurons in each domain in different modules of MLLMs. Bold is used to highlight the domain with the most neurons in each module.
We identify domain-specific neurons using the method described in Section 3.2. Since neurons in different modules may have different activation patterns, as shown in Appendix C, we detected those domain-specific neurons module by module. Figure 5 shows the distribution of domain-specific neurons for each layer in each module of MLLMs.
Three-stage framework of LLM understanding multimodal features.
Two obvious turning points can be observed in both LLaVA-NeXT and InstructBLIP’s language model, one in the intermediate layer and the other near the output layer. Inspired by Zhao et al. (2024c), we thus propose a three-stage framework of LLM understanding multimodal features: 1) In the first several layers, projected features are further aligned with word space. Around the turning point, the multimodal features are embedded into a uniform representation space, where included domain-specific information needs to be processed by more domain-specific neurons. 2) Transitioning into the second phase, features are further generalized and understood by language models, where domain-specific neurons decrease sharply. 3)In the third stage, language models generate responses to the input, showing a rise of neurons specific to target tasks.
Our hypothesis aligns with the previous conclusion on smaller multimodal models like GPT-J Wang and Komatsuzaki (2021), as Schwettmann et al. (2023) argue that outputs of the projection layer are further translated in deeper layers within the transformer. To further validate our hypothesis, we employ logit lens to visualize the transformation of multimodal features within language models in Section 4.2.3.
Domain-specific information in different semantic levels.
Domain-specific neurons are mainly distributed in shallow and intermediate layers within MLLMs’ vision encoders. Prior research discussed the correlation between the semantic level and layer depth, which found that more deep layers will focus on higher-level concepts in visual networks. In our settings, the document domain contains more low-level concepts, such as line and shape, while the remote sensing and medical domain may include more high-level concepts, like architectures and organs. Therefore, document neurons are mainly gathered in bottom layers close to the input end. Another interesting phenomenon is the rise of auto driving neurons near the output layer of InstructBLIP’s Q-Former, we conjecture this may reflect the struggle of model to understand the language instructions of auto driving domain.
Gap between the ability of MLLM to handle visual and lingual instructions.
Table 1 demonstrates the number of neurons in each domain. Remote sensing neurons have the largest proportion in LLaVA-NeXT’s vision encoder, MLP projector and language model, while in InstructBLIP, the domain owns most specific neurons are document, auto driving and auto driving separately. We argue that the number of specific neurons reflects the understanding ability of MLLM in the target domain, as more specific neurons may mean more demanding to process domain-specific information. In contrast, less specific neurons mean more generalized features in the target domain Tang et al. (2024). In this way, we find that there exists a large visual gap between domains like remote sensing, document and medical, comparing the two domains left. Moreover, InstructBLIP seems less proficient in processing questions from auto driving, as neurons of this domain exhibit the highest number in Q-Former and LLM. There is also a similar pattern in its language model as for the auto driving domain. In other words, while visual features of auto driving domain can be processed well by existing vision encoder, the language instruction of this domain may be hard to handle for language model.
4.2.2 Influence of domain-specific neurons
Perturbation for Performance in VQA Tasks
Model Deactivated Module(s) VQAv2 PMC-VQA LingoQA RS-VQA DocVQA
LLaVA-NeXT None 74.9 34.4 20.6 42.5 59.2
Vision Encoder 75.8 34.3 24.6 42.1 58.3
MLP Projector 74.9 34.4 24.2 42.5 59.2
LLM 75.7 34.5 24.2 41.0 59.0
All 73.5 34.5 24.2 38.5 57.0
InstructBLIP None 66.1 28.1 20.6 34.7 24.0
Vision Encoder 66.9 31.0 21.8 34.8 23.8
Q-Former 67.1 32.4 20.0 33.1 24.6
LLM 67.1 32.6 24.2 35.5 24.4
All 68.6 30.9 18.0 33.6 23.8
Table 2: Accuracy (%) of LLaVA-NeXT and InstructBLIP on selected domains with corresponding domain-specific neurons deactivated. “None” means no neurons are deactivated, while “All” means deactivating domain-specific neurons in all the modules above. Bold is used to highlight the worst performance in each column.
Table 2 demonstrates the performance of LLaVA-NeXT and InstructBLIP after deactivating domain-specific neurons in different modules. While the performance decrease after deactivating is slight for most domains, we find that deactivating remote sensing neurons in LLaVA-NeXT and auto driving neurons in InstructBLIP will result in a great fall of 4.0 and 2.6 accuracy separately. Similarly, in the document domain, deactivating domain-specific neurons at most causes a 2.2 accuracy decrease for LLaVA-NeXT. Interestingly, in some cases, removing domain-specific information seems to benefit the target task, as the accuracy of LLaVA-NeXT in auto driving has risen from 20.6 to 24.2. We leave this for future work.
In summary, deactivating domain-specific neurons will not cause a sharp decrease in performance for some domains. To investigate the reason for that further, we compare the influence of domain-specific neurons in MLLMs’ hidden states.
Perturbation for Hidden States
Baseline Module VQAv2 PMC-VQA LingoQA DocVQA RS-VQA
LLaVA-NeXT Random (Avg.) 8.41 18.90 16.04 21.81 32.76
LLM 0.01 0.01 0.02 0.10 0.02
Vision Encoder 17.19 30.98 35.74 46.75 49.90
MLP Projector 0.0 0.0 0.0 0.0 0.0
All 17.19 30.98 35.74 46.75 49.90
InstructBLIP Random (Avg.) 5.13 8.15 8.57 14.85 9.91
LLM 6.84 12.13 9.62 7.80 11.98
Vision Encoder 2.44 17.93 5.33 26.11 23.76
Q-Former 2.93 11.61 6.95 14.58 6.52
All 8.00 24.84 12.77 29.04 26.58
Table 3: The deviation (%) of hidden states of MLLMs’ last layer after deactivating domain-specific neurons. We calculate the deviation $d$=$\frac{\|H_{n}-H_{d}\|_{2}}{\|H_{n}\|_{2}}$, where $H_{n}$ and $H_{d}$ denotes the hidden states before and after deactivating neurons separately. Bold is used to highlight the largest deviation in each column. Random (Avg.) refers to the average deviation by randomly deactivating neurons of the same number in all modules.
We demonstrate the influence of domain-specific neurons on MLLM’s last hidden states in Table 3. Surprisingly, deactivating domain-specific neurons causes a large perturbation to LLaVA-NeXT and InstructBLIP’s hidden states. In contrast, deactivating all of the domain-specific neurons can have little effect on the accuracy of these domains, as shown in Table 2. Therefore, we argue that both LLaVA and InstructBLIP fail to take full advantage of the domain-specific information in specific domains, which may limit their cross-domain ability. In other words, the representations within MLLM’s language models are highly generalized.
4.2.3 Case Study
(a) Visaul and language input. The area in the image is located in New York.
(b) The next token distribution of the second image token, the expected next token is ‘</s>’ (i.e., end of sentence).
(c) The next token distribution of the last text token, the expected next token is the correct answer ‘no’.
Figure 6: The logit lens can be applied to decode the hidden states of the language model’s intermediate layers into the probability distribution of the vocabulary. We only display the top 5 candidates for each layer in the heatmap. Color indicates the probability of candidates from low (white) to high (blue).
(d) Average entropy of next-token distribution of InstructBLIP.
(e) Average entropy of next-token distribution of LLaVA-NeXT.
Figure 7: The average entropy of next token probability distribution for image and text tokens. The colors of lines denote different domains, such as auto driving (ad), remote sensing (rs), medical (med), common (com), and document (doc). We use dashed lines and solid lines to distinguish curves of image and text tokens.
To investigate how MLLM’s language model processes image tokens, we employ logit lens nostalgebraist (2020) to decode the hidden states of the language model’s intermediate layers into the probability of the next token across the vocabulary. As displayed in Figure 6, when feeding a remote sensing image-question pair into InstructBLIP, we get that the most likely token next to the second image token is ’</s>’, while the most likely token next of the last text token is the correct answer, ’no’. Interestingly, two place names, "Hermann" and "Baltimore", have appeared among the top token candidates when the image input is a remote sensing picture of New York. In multilingual literature, similar phenomena have also been observed. For instance, when Llama 2 receives the French token ’fleur’ in the input, the English concept ’__flower’ will appear in the intermediate distribution Wendler et al. (2024). This suggests that the decoded vocabulary distribution can to some extent reflect the semantic concepts understood by the language model. Despite this observation, we note that the decoded distribution of image tokens is far more sparse than text tokens; even in the output layer, the probability of the most likely next token ’</s>’ is lower than 40%. It indicates that projected tokens may be treated as a sparse mixture of concepts in the representation space instead of a simple word. We also demonstrate more cases of logit lens in different domains in Appendix D.
To further explore this phenomenon, we calculate the average entropy of the next token distribution for image tokens and text tokens separately, as shown in Figure 7. As the curves of image tokens tend to be above those of text tokens for all the layers, the next token distributions of image tokens are indeed more sparse than those of text tokens. Moreover, the tendency of entropy curves aligns with the hypothesis we have proposed in Section 4.2.1. In the first stage, features are aligned into a uniform representation space, where entropy curves level off high. In the second stage, the language model understands and processes the information, as curves drop sharply in the intermediate layers. Finally, the model selects the suitable next token to output, resulting in a slight increase in entropy. A similar tendency has also been observed in English-native multilingual LLMs when handling non-English inputs Wendler et al. (2024).
5 Conclusion
To explore the neuron-level domain-specific interpretation in current MLLMs, we propose MMNeuron framework inspired by multilingual research. In particular, we first calculate the activation probabilities of neurons in LLaVA-NeXT and InstructBLIP across five domains, identifying those with low domain DAPE scores as domain-specific neurons. By analyzing the distribution of domain-specific neurons and their influence on MLLMs, we find that the language model modules of MLLMs fail to fully utilize domain-specific information in VQA tasks. We further propose a three-stage framework that the language model module employs to handle projected visual features and corroborate it indirectly with logit lens. We envision that our work will shed light on the interpretability of current MLLMs, aiding the development of cross-domain, all-encompassing MLLMs in the future.
6 Limitations
Despite the findings we demonstrate in our work, there still exist several limitations:
  • 1.
    Our experiments are conducted mainly on LLaVA-NeXT and InstructBLIP, whose frameworks are similar in aligning visual features with the word embedding space via a projector. This means that our findings may not be directly applicable to models that utilize different frameworks, such as those injecting vision representations into LLMs by layer Wemm (2023).
  • 2.
    Although we find that domain-specific information is not fully utilized by the language model modules of MLLMs, how such information is conveyed and ignored between different layers is still less known. We leave these problems for future work.
  • 3.
    We discuss the possible workflow of the language model module handling projected visual features through logit lens. While there do exist special semantic concepts in the decoded representations, we still know little about how these concepts are encoded and how projected features interact with word embeddings during the forward pass. Therefore, further mathematical analysis in this area is still required in the future.
Appendix A Visual Domain Definition
Domain Definition Dataset Num of Samples Example
Common Scenes Natural images taken in everyday life VQAv2 Goyal et al. (2017) 21K [Uncaptioned image]
Remote Sensing Images captured by remote sensing sensors such as satellites RS-VQA Lobry et al. (2020) 11K [Uncaptioned image]
Medical Medical images obtained through techniques like CT and X-ray PMC-VQA Zhang et al. (2023b) 15K [Uncaptioned image]
Document Documents containing charts, text-rich images, and records DocVQA Mathew et al. (2021) 10K [Uncaptioned image]
Auto Driving Scenes captured from the viewpoint of a vehicle’s camera LingoQA Marcu et al. (2023) 14K [Uncaptioned image]
Table 4: Domain definition and the corresponding datasets.
We define five domains in this work and each of them has characterized image features, as displayed in Table 4.
Appendix B Prompt Template
B.1 Instructions templates for VQA
Step Model Prompt
Identification LLaVA-NeXT <Image><Question>
InstructBLIP <Image><Question>
Evaluation (open-ended) LLaVA-NeXT A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. USER: <Image> {Role Description}* Question: {Question} Context: N/A Answer the question using a single word or phrase. ASSISTANT:
InstructBLIP <Image> {Role Description}* Question: {Question} Short Answer:
Evaluation (multi-option) LLaVA-NeXT A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. USER: <Image> Question: {Question} Context: N/A Options: {Options} Answer with the option’s letter from the given choices directly. ASSISTANT:
InstructBLIP <Image> Question: {Question} Options: {Options} Answer with the option’s letter from the given choices directly.
Table 5: Prompt templates we have used in different steps. For identifying domain-specific neurons, plain questions are input into models. During evaluation, we follow the templates provided by official repositories or codes.
For instructions with options, we separate options in alphabetical order, as shown in Appendix B.2. $\star$ : A role description has been provided to help models better understand the tasks in auto driving. As shown below:
“Role: You are an advanced AI assistant installed on the Ego vehicle, equipped with conversational analysis capabilities for discussing autonomous driving scenarios. The perspective presented is from the point-of-view of the Ego vehicle, where the camera is mounted. It’s important to note that the Ego vehicle itself is not visible in the images provided.”
B.2 Prompt Examples
(a) Prompt example for open-ended tasks, the image and question come from RSVQA.
(b) Prompt example for multi-option tasks, the image and question come from PMC-VQA.
Figure 8: Prompt examples of conversational format for LLaVA-NeXT.
We display the prompt format we use for evaluation in LLaVA-NeXT, as shown in Figure 8. The prompt for InstructBLIP come from direct format in Table B.1.
Appendix C Silent Neurons in MLLM’s Vision Encoder
(a) Ratio of silent and activated neurons in IntructBLIP’s vision encoder.
(b) Ratio of silent and activated neurons in LLaVA-NeXT’s vision encoder.
Figure 9: Layer-wise distribution of silent neurons.
We observed that several neurons in the vision encoders of LLaVA-NeXT and InstructBLIP remain silent regardless of the input images. We refer to these as “silent neurons". Figure 9 illustrates the distribution of these silent neurons within the vision encoders.
Appendix D Logit Lens Cases
We provide more cases from other four datasets, as displayed in Figure 10, 11, 12 and 13. For LingoQA (auto driving domain), the visual inputs for each question are multiple images.
(a) Visual and language input of PMC-VQA.
(b) The next token distribution of the 8th image token.
(c) The next token distribution of the last text token.
Figure 10: Case of logit lens in InstructBLIP on PMC-VQA.
(a) Visual and language input of DocVQA.
(b) The next token distribution of the 377th image token.
(c) The next token distribution of the 5th from last text token.
Figure 11: Case of logit lens in LLaVA-NeXT on DocVQA.
(a) Visual and language input of VQAv2.
(b) The next token distribution of the 49th image token.
(c) The next token distribution of the 9th from last text token.
Figure 12: Case of logit lens in LLaVA-NeXT on VQAv2.
(a) Images inputs of LingoQA. Question: Is there a vehicle ahead of you in your lane?
(b) The next token distribution of the 37th image token in LLaVA-NeXT’s vision encoder.
(c) The next token distribution of the 18th from the last text token.
Figure 13: Case of logit lens in LLaVA-NeXT on LingoQA.